All the real zeros of the given polynomial are integers. Find the zeros.
(Enter your answers as a comma-separated list. Enter all answers including repetitions.)
P(x) = x^4 − 10x^2 + 9
x =
Write the polynomial in factored form.
P(x) =

Answer by ikleyn(52847)

.
All the real zeros of the given polynomial are integers. Find the zeros.
(Enter your answers as a comma-separated list. Enter all answers including repetitions.)
P(x) = x^4 − 10x^2 + 9
x =
Write the polynomial in factored form.
P(x) =
~~~~~~~~~~~~~~~


Make the decomposition / (factoring) in two steps


(1)  x^4 - 10x^2 + 9 = (x^2-9)*(x^2-1)   (as if you factor mentally a quadratic polynomial).


(2)  Next, factor each of the participating quadratic polynomials separately

           x^2 - 10x^2 + 9 = (x^2-9)*(x^2-1) = (x+3)*(x-3)*(x+1)*(x-1).

Done.